OpenStudy (anonymous):

Subtract:(3x2 - x - 5) - (x + 2)

OpenStudy (anonymous):

is it 3 times 2 or 3 times x2

OpenStudy (anonymous):

x2

OpenStudy (anonymous):

ok then (6x- x-5)-(x+2) =5x-5-x+2 =5x-x-5+2 =4x-3